OpenStudy (anonymous):

which of the following represents (going to write the equation) in exponential form?

OpenStudy (anonymous):

\[^4\sqrt{x^3}\]

OpenStudy (anonymous):

@charlotte123

OpenStudy (charlotte123):

so its 4 sqrt x^2? :)

OpenStudy (anonymous):

4/3? @charlotte123

OpenStudy (charlotte123):

can u write it out for me? XD

OpenStudy (anonymous):



OpenStudy (anonymous):

@charlotte123

OpenStudy (charlotte123):

can u try to write it in this box? the writing one XD I still can't get what it says >.> is that x 4/3 or r u just trying to find 4/3?

OpenStudy (anonymous):

lol its x4/3 but im not sure if its that or x3/4

OpenStudy (anonymous):

@charlotte123

OpenStudy (charlotte123):

if it was x4/3, x would = 0 :) if it was x3/4 x would = 0 :D

OpenStudy (anonymous):

r u saying its neither??? @charlotte123

OpenStudy (anonymous):